# Unlocking Textual and Visual Wisdom: Open-Vocabulary 3D Object Detection Enhanced by Comprehensive Guidance from Text and Image

### v1

@@ Abstract
Open-vocabulary 3D object detection (OV-3DDet) aims to localize and recognize both seen and previously unseen object categories within any new 3D scene. While language and vision foundation models have achieved success in handling various open-vocabulary tasks with abundant training data, OV-3DDet faces a significant challenge due to the limited availability of training data. Although some pioneering efforts have integrated vision-language models (VLM) knowledge into OV-3DDet learning, the full potential of these foundational models has yet to be fully exploited. In this paper, we unlock the textual and visual wisdom to tackle the open-vocabulary 3D detection task by leveraging the language and vision foundation models. We leverage a vision foundation model to provide image-wise guidance for discovering novel classes in 3D scenes. Specifically, we utilize a object detection vision foundation model to enable the zero-shot discovery of objects in images, which serves as the initial seeds and filtering guidance to identify novel 3D objects. Additionally, to align the 3D space with the powerful vision-language space, we introduce a hierarchical alignment approach, where the 3D feature space is aligned with the vision-language feature space using a pre-trained VLM at the instance, category, and scene levels. Through extensive experimentation, we demonstrate significant improvements in accuracy and generalization, highlighting the potential of foundation models in advancing open-vocabulary 3D object detection in real-world scenarios.
@@ 1 Introduction
3D object detection serves as a fundamental component in understanding 3D scenes, playing a pivotal role in various applications [1, 13, 21, 5] such as autonomous driving and robot interaction. However, conventional approaches to 3D object detection [28, 14, 24, 22, 15] often operate under the assumption that the detection targets during testing remain consistent with those observed during training. This assumption fails to reflect the dynamic and evolving nature of real-world scenarios, where the objects within scenes can vary and expand over time. Consequently, the capability of open-vocabulary 3D object detection, which enables the localization and recognition of both seen and previously unseen objects within new scenes, becomes essential for their practical deployment in real-world settings.
@@ 2 Related Work
3D Object Detection endeavors to locate and identify 3D objects within a given scene, with numerous approaches proposed to address this challenge. VoteNet [16] introduced a point voting strategy, leveraging PointNet [17] for processing 3D points. Evolving from VoteNet, MLCVNet [19] introduces additional modules for point voting, aiming to capture multi-level contextual information and enhance overall detection performance. The advent of transformer-based methods has significantly shaped the landscape of 3D object detection. Notably, GroupFree [11] utilizes a transformer as the prediction head, eliminating the need for manually crafted grouping and harnessing the transformer’s capabilities to improve detection accuracy. In recent years, the trend has shifted towards end-to-end models, with 3DETR [14] standing out as the pioneer in employing an end-to-end transformer architecture for 3D object detection, utilizes bipartite graph matching to establish associations between predictions and ground truth. However, traditional 3D object detection methods only detect objects seen during the training stage and can’t handle the open-vocabulary scenario.
@@
Open-vocabulary Object Detection aims to detect objects that include both seen and unseen classes in an image. Since the success of CLIP in associating language and the 2D domain, research on zero-shot learning and open-vocabulary has become a trend. For example, Detic [27] uses ImageNet to train the classifiers of detectors, expanding the vocabulary of detectors to tens of thousands of concepts. GLIP [8] and MDETR [6] treat detection as the grounding task and adopt a text query for the input image to predict corresponding boxes. GLIPv2 [25] reformulates the detection task, introduces a novel region-word level contrastive learning task, and includes masked language modeling. The success of open-vocabulary 2D detection methods provides the potential to facilitate OV-3DDet. Our method also leverages a pre-trained language-vision model to provide guidance for 3D detection.
@@
Open-vocabulary 3D Object Detection (OV-3DDet) presents a substantial challenge, aiming to detect and locate objects in a 3D space, encompassing both known and unknown object categories. In OV3DET [12], a 2D detector is employed to generate pseudo-labels. These labels are utilized for training a class-agnostic detector, followed by a contrastive learning step that aligns the 3D feature space with image and language features. L3DET [29] contributes to the field by enriching 3D scene datasets through the introduction of novel objects and associated text descriptions. It leverages cross-domain, category-level contrastive learning to align feature spaces between point clouds and language, facilitating effective cross-modal reasoning. CoDA [2] takes a distinctive approach by integrating 3D box geometry priors and 2D semantic open-vocabulary priors for novel object identification. Discovered novel objects contribute to subsequent detector training, aligning the 3D feature space with the vision-language feature space through class-agnostic knowledge distillation and class-aware contrastive learning. However, CoDA exhibits inaccuracies and insufficiencies in novel class discovery and feature alignment. While PLA [4] introduces alignment at the instance, category, and scene levels, its primary focus lies within the segmentation setting, resulting in a higher cost at the point-wise level rather than the object-wise level. In response to these challenges, we propose an alternative approach.
@@ 3.2 INHA Overview
Our proposed Image-guided Novel class discovery and Hierarchical feature space Alignment (INHA) approach adopts 3DETR [14] as our 3D object detector, which comprises PointNet+++++ +, Transformer encoder, Transformer decoder, a bounding box regression head, and a classification head. The training process consists of three stages, illustrated in Figure 2. In the first stage, we train a base class-agnostic 3D object detector. More specifically, we remove the classification-related training and solely utilize bounding box regression loss $\mathcal{L}_{\text{box}}$ [14] for training the detector. After the first stage, the detector is capable of detecting objects in a point cloud without considering their class information. In the second stage, we incorporate both the 2D detector and the 3D detector to discover novel objects (cf. Sec. 3.3). The discovered objects are stored and then used for further training the 3D detector. This process enhances the 3D detector’s ability to handle novel classes. In the final stage, we introduce a hierarchical paradigm (cf. Sec. 3.4) that aligns the feature space at instance, class, and scene levels.
@@ 3.3 Image-Guided Novel Class Discovery
The image provides rich appearance cues that aid in identifying cluttered objects. Leveraging the capabilities of recent open-vocabulary 2D object detectors, even extremely small or occluded novel objects can be identified in images. For instance, a distant object can be easily discerned by a 2D detector using only a few pixels. In contrast, 3D object detectors struggle to recognize distant objects with very few points. However, 3D object detectors excel in capturing rich 3D geometrical information, resulting in more precise localization predictions. Recognizing the complementary nature of these modalities, we present the Image-Guided Novel Class Discovery (IGND) module, illustrated in Figure 3, to discover more 3D novel objects. In this module, we utilize a pre-trained open-vocabulary 2D detector, i.e. Detic [27], to extract valuable object-level information (2D object bboxes) from images. This information is then effectively integrated with valuable 3D data to guide the discovery of 3D novel objects. Specifically, this guidance occurs in two key steps: a) lifting the centroids of the 2D objects to 3D space to provide supplementary query seeds for generating more 3D object proposals, and b) utilizing the bounding boxes of the 2D objects to select reliable novel 3D bounding boxes. Through these two guided processes, we enhance the recall rates of novel classes, as depicted in Figure 6.
@@
$\hat{\textbf{b}}^{\mathcal{N}}=\left\{\hat{\textbf{b}}_{i}|\eta(\textbf{b}_{\text{m}},\hat{\textbf{b}}_{i})\geq\max_{\eta_{mj}}\eta(\textbf{b}_{m},\hat{\textbf{b}}_{j}),\eta_{mi}\geq\epsilon,\textbf{c}_{m}\in C^{\mathcal{N}}\right\}_{m=1}^{M},$ (1)
where $\epsilon$ is a threshold used to filter out cases with low IoU. We then match the corresponding 3D objects based on these selected 2D boxes. These 3D boxes are stored in the novel object memory bank and are subsequently used to retrain the 3D detector along with the base objects. Periodically, we conduct the discovery procedure and clear the previous memory bank to remove inaccurate objects.
@@ 3.4 Hierarchical Cross-modal Feature Alignment
The alignment at the instance level emphasizes the consistency in 3D features and image features. While not considering general class information in the language domain yet.
@@
$\begin{split}\mathcal{L}_{cls}^{3d\Rightarrow rgb,text}=-\frac{1}{S}\sum_{i=1}^{S}\log\frac{{\textstyle\sum_{k=1}^{S}}\mathds{1}(i\neq k,c_{i}=c_{k})\mathrm{e}^{\textbf{g}_{i}\cdot\textbf{g}_{k}/\tau}}{\sum_{j=1}^{S}\mathrm{e}^{\textbf{g}_{i}\cdot\textbf{g}_{j}/\tau}}.\end{split}$ (3)
Here, $\tau$ is the temperature parameter, $\mathds{1}(\cdot)$ is the indicator function, which yields 1 when the condition is met and 0 otherwise.
Scene-level Cross-modal Alignment. The objects within a scene often exhibit strong correlations, where certain objects are more likely to coexist. For example, in a living room, a bed is typically accompanied by a dresser, but not a refrigerator. Utilizing this correlation prior is beneficial for aligning the 3D scene-level feature space. Here, we align the 3D scene features with the language semantic feature space at the scene level. To generate the text scene feature, we first create a scene-level caption containing all class names present in the scene. This caption is then processed by the CLIP text encoder to produce the scene-level text feature $\left\{\textbf{z}_{i}^{text}\right\}_{i=1}^{L}$, $L$ signifies the number of scenes in a batch. For scene-level 3D feature extraction, we introduce the Permutation-Invariant Scene-level feature Extraction (PISE) module. This involves concatenating all individual 3D object features within a scene and projecting them into a high-dimensional space. To address permutation invariance, we utilize a max pooling operation on these high-dimensional features. The detailed architecture of the PISE module is illustrated in Figure 5. Let $\left\{\textbf{z}_{i}^{3D}\right\}_{i=1}^{L}$ represent the scene-level 3D feature extracted by the PISE module. Subsequently, we employ a contrastive loss to align the 3D scene feature with the corresponding text scene feature:
$\mathcal{L}_{scene}^{3d\Rightarrow text}=-\frac{1}{L}\sum_{i=1}^{L}\log\frac{\mathrm{e}^{\textbf{z}^{3d}_{i}\cdot\textbf{z}^{text}_{i}}}{\sum_{j=1}^{L}\mathrm{e}^{\textbf{z}^{3d}_{i}\cdot\textbf{z}^{text}_{j}}}.$ (4)
@@ 4.1 Datasets
Our proposed approach is evaluated on two challenging 3D indoor detection datasets: SUN RGB-D, which consists of 5,285 training samples with oriented 3D bounding box labels for 46 object categories. We select the 10 most frequent classes as base classes and the remaining 36 most frequent classes as novel classes; ScanNetv2 [3], it has 1,201 training samples with 200 object categories. We use the 10 most frequent classes as base classes and the remaining 50 most frequent classes as novel classes. All configurations above are the same as in [2].
@@ 4.3 Implementation Details
Training Strategy and Hyperparameters. Our training procedure includes 3 stages. In the first stage, i.e. base detector training, we train the 3D detector for 1000 epochs for a fair comparison with [2]. In the second stage, i.e. novel class discovery, we train for an additional 200 epochs. In the final stage, i.e. feature space alignment, we train another 50 epochs to align the 3D feature space to the vision-language feature space. For the first stage, we configure a batch size of 8 and a learning rate of 0.0004, utilizing 128 queries. Subsequently, during the second and third stages, we specifically adjust the learning rate to 0.0001, batch size to 16, and increase the query size to 196 for enhanced detector training and feature alignment. We set the temperature $\tau$ to 1.0 and the threshold $\epsilon$ to 0.75. Additionally, the hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are initially set to 0.02 during warm-up and later adjusted to 1, 1, and 0.5, respectively.
Model Selection and Prompt Setting. We modify 3DETR [14] as our 3D detector, removing the classification head and using only the bounding box regression loss. Additionally, we employ the pretrained Detic [27] as our 2D detector. For feature alignment, we leverage the pretrained CLIP [20] to encode image and text features. In the class-level and scene-level feature space alignment, we generate text prompts for language feature encoding. Specifically, for the class-level prompt, we generate feature embeddings using the template “A photo of [class name].”, where [class name] represents the category name. For the scene-level prompt, we concatenate the class names in a scene into a list and insert the list into the template “A room with [class list].”, where [class list] denotes the position to insert the list. Additionally, we include the non-object description prompt, i.e., “A photo of nothing.” to represent areas without discriminative objects.
Table 1: Comparison results(%) of our INHA and baseline methods on SUN RGB-D.
@@ 4.4 Main Results
Method toile bed chair sofa dresser table cabinet bookshelf pillow sink Avg.
L3DET [29] 56.34 36.15 16.12 23.02 8.13 23.12 14.73 17.27 23.44 27.94 24.626
OV-3DET [12] 57.29 42.26 27.06 31.5 8.21 14.17 2.98 5.56 23 31.6 24.363
CoDA [2] 68.09 44.04 28.72 44.57 3.41 20.23 5.32 0.03 27.95 45.26 28.762

### v2

@@ Abstract
Open-vocabulary 3D object detection (OV-3DDet) aims to localize and recognize both seen and previously unseen object categories within any new 3D scene. While language and vision foundation models have achieved success in handling various open-vocabulary tasks with abundant training data, OV-3DDet faces a significant challenge due to the limited availability of training data. Although some pioneering efforts have integrated vision-language models (VLM) knowledge into OV-3DDet learning, the full potential of these foundational models has yet to be fully exploited. In this paper, we unlock the textual and visual wisdom to tackle the open-vocabulary 3D detection task by leveraging the language and vision foundation models. We leverage a vision foundation model to provide image-wise guidance for discovering novel classes in 3D scenes. Specifically, we utilize a object detection vision foundation model to enable the zero-shot discovery of objects in images, which serves as the initial seeds and filtering guidance to identify novel 3D objects. Additionally, to align the 3D space with the powerful vision-language space, we introduce a hierarchical alignment approach, where the 3D feature space is aligned with the vision-language feature space using a pretrained VLM at the instance, category, and scene levels. Through extensive experimentation, we demonstrate significant improvements in accuracy and generalization, highlighting the potential of foundation models in advancing open-vocabulary 3D object detection in real-world scenarios.
@@ 1 Introduction
3D object detection serves as a fundamental component in understanding 3D scenes, playing a pivotal role in various applications [1, 14, 22, 6] such as autonomous driving and robot interaction. However, conventional approaches to 3D object detection [30, 15, 25, 23, 16, 28, 21, 5] often operate under the assumption that the detection targets during testing remain consistent with those observed during training. This assumption fails to reflect the dynamic and evolving nature of real-world scenarios, where the objects within scenes can vary and expand over time. Consequently, the capability of open-vocabulary 3D object detection, which enables the localization and recognition of both seen and previously unseen objects within new scenes, becomes essential for their practical deployment in real-world settings.
@@ 2 Related Work
3D Object Detection endeavors to locate and identify 3D objects within a given scene, with numerous approaches proposed to address this challenge. VoteNet [17] introduced a point voting strategy, leveraging PointNet+++++ +[18] for processing 3D points. Evolving from VoteNet, MLCVNet [19] introduces additional modules for point voting, aiming to capture multi-level contextual information and enhance overall detection performance. The advent of transformer-based methods has significantly shaped the landscape of 3D object detection. Notably, GroupFree [12] utilizes a transformer as the prediction head, eliminating the need for manually crafted grouping and harnessing the transformer’s capabilities to improve detection accuracy. Recently, the trend has shifted towards end-to-end models, with 3DETR [15] standing out as the pioneer in employing an end-to-end transformer architecture for 3D object detection. It utilizes bipartite graph matching to establish associations between predictions and ground truths. However, traditional 3D object detection methods only detect objects seen during the training stage and cannot handle the open-vocabulary scenario.
@@
Open-vocabulary Object Detection aims to detect objects that include both seen and unseen classes in an image. Since the success of CLIP in associating language and the 2D domain, research on zero-shot learning and open-vocabulary has become a trend. For example, Detic [29] uses ImageNet to train the classifiers of detectors, expanding the vocabulary of detectors to tens of thousands of concepts. GLIP [9] and MDETR [7] treat detection as the grounding task and adopt a text query for the input image to predict corresponding boxes. GLIPv2 [26] reformulates the detection task, introduces a novel region-word level contrastive learning task, and includes masked language modeling. The success of open-vocabulary 2D detection methods provides the potential to facilitate OV-3DDet. Our method also leverages a pretrained language-vision model to provide guidance for 3D object detection.
@@
Open-vocabulary 3D Object Detection (OV-3DDet) presents a substantial challenge, aiming to detect and locate objects in a 3D space, encompassing both known and unknown object categories. In OV3DET [13], a 2D detector is employed to generate pseudo-labels. These labels are utilized for training a class-agnostic detector, followed by a contrastive learning step that aligns the 3D feature space with image and language features. L3DET [31] contributes to the field by enriching 3D scene datasets through the introduction of novel objects and associated text descriptions. It leverages cross-domain, category-level contrastive learning to align feature spaces between point clouds and language, facilitating effective cross-modal reasoning. CoDA [2] takes a distinctive approach by integrating 3D box geometry priors and 2D semantic open-vocabulary priors for novel object discovery. Discovered novel objects contribute to subsequent detector training, aligning the 3D feature space with the vision-language feature space through class-agnostic knowledge distillation and class-aware contrastive learning. However, CoDA exhibits inaccuracies and insufficiencies in novel class discovery and feature alignment. While PLA [4] introduces alignment at the instance, category, and scene levels, its primary focus lies within the segmentation setting, resulting in a higher cost at the point-wise level rather than the object-wise level. In response to these challenges, we propose an alternative approach.
@@ 3.2 INHA Overview
Our proposed Image-guided Novel class discovery and Hierarchical feature space Alignment (INHA) approach adopts 3DETR [15] as our 3D object detector, which comprises PointNet+++++ +, Transformer encoder, Transformer decoder, a bounding box regression head, and a classification head. The training process consists of three stages, illustrated in Figure 2. In the first stage, we train a base class-agnostic 3D object detector. More specifically, we remove the classification-related training and solely utilize bounding box regression loss $\mathcal{L}_{\text{box}}$ [15] for training the detector. After the first stage, the detector is capable of detecting objects in a point cloud without considering their class information. In the second stage, we incorporate both the 2D detector and the 3D detector to discover novel objects (cf. Sec. 3.3). The discovered objects are stored and then used for re-training the 3D detector. This process enhances the 3D detector’s ability to handle novel classes. In the final stage, we introduce a hierarchical paradigm (cf. Sec. 3.4) that aligns the feature space at instance, class, and scene levels.
@@ 3.3 Image-Guided Novel Class Discovery
The image provides rich appearance cues that aid in identifying cluttered objects. Leveraging the capabilities of recent open-vocabulary 2D object detectors, even extremely small or occluded novel objects can be identified in images. For instance, a distant object can be easily discerned by a 2D detector using only a few pixels. In contrast, 3D object detectors struggle to recognize distant objects with very few points. However, 3D object detectors excel in capturing rich 3D geometrical information, resulting in more precise localization predictions. Recognizing the complementary nature of these modalities, we present the Image-Guided Novel Class Discovery (IGND) module, illustrated in Figure 3, to discover more 3D novel objects. In this module, we utilize a pretrained open-vocabulary 2D detector, i.e. Detic [29], to extract valuable object-level information (2D object bboxes) from images. This information is then effectively integrated with valuable 3D data to guide the discovery of 3D novel objects. Specifically, this guidance occurs in two key steps: a) lifting the centroids of the 2D objects to 3D space to provide supplementary query seeds for generating more 3D object proposals, and b) utilizing the bounding boxes of the 2D objects to select reliable novel 3D bounding boxes. Through these two guided processes, we enhance the recall rates of novel classes, as depicted in Figure 5.
@@
$\hat{\textbf{b}}^{\mathcal{N}}=\left\{\hat{\textbf{b}}_{i}|\eta(\textbf{b}_{\text{m}},\hat{\textbf{b}}_{i})\geq\max_{\eta_{mj}}\eta(\textbf{b}_{m},\hat{\textbf{b}}_{j}),\eta_{mi}\geq\epsilon,c_{m}\in C^{\mathcal{N}}\right\}_{m=1}^{M},$ (1)
where $\epsilon$ is a threshold used to filter out cases with low IoU. We take the corresponding 3D objects from these selected 2D boxes $\hat{\textbf{b}}^{\mathcal{N}}$ and store them in the novel object memory bank. These novel 3D objects are subsequently used to retrain the 3D detector along with the base objects. Periodically, we conduct the discovery procedure and update the entire memory bank with the newly discovered novel objects.
@@ 3.4 Hierarchical Cross-modal Feature Alignment
The alignment at the instance level emphasizes the consistency in 3D features and image features, without yet considering general class information in the language domain.
@@
$\begin{split}\mathcal{L}_{cls}^{3d\Rightarrow rgb,text}=-\frac{1}{S}\sum_{i=1}^{S}\log\frac{{\textstyle\sum_{k=1}^{S}}\mathds{1}(i\neq k,c_{i}=c_{k})\mathrm{e}^{\textbf{g}_{i}\cdot\textbf{g}_{k}/\tau_{1}}}{\sum_{j=1}^{S}\mathrm{e}^{\textbf{g}_{i}\cdot\textbf{g}_{j}/\tau_{1}}}.\end{split}$ (3)
Here, $\tau_{1}$ is the temperature parameter, $\mathds{1}(\cdot)$ is the indicator function, which yields 1 when the condition is met and 0 otherwise.
Scene-level Cross-modal Alignment. The objects within a scene often exhibit strong correlations, where certain objects are more likely to coexist. For example, in a living room, a bed is typically accompanied by a dresser, but not a refrigerator. Utilizing this correlation prior is beneficial for aligning the cross-modal feature spaces at the scene level. Here, we align the 3D scene features with the text scene features. To generate the text scene feature, we first create a scene-level caption containing all class names present in the scene. This caption is then processed by the CLIP text encoder to produce the scene-level text feature $\textbf{z}^{text}$. For scene-level 3D feature extraction, we introduce the Permutation-Invariant Scene-level feature Extraction (PISE) module. This involves concatenating all individual 3D object features within a scene and projecting them into a high-dimensional space. To address permutation invariance, we utilize a max pooling operation on these high-dimensional features. The detailed architecture of the PISE module is illustrated in Figure 4. With the scene-level 3D feature $\textbf{z}^{3D}$ extracted by the PISE module, we employ a contrastive loss to align the 3D scene features with the corresponding text scene features within a batch:
$\mathcal{L}_{scene}^{3d\Rightarrow text}=-\frac{1}{L}\sum_{i=1}^{L}\log\frac{\mathrm{e}^{\textbf{z}^{3d}_{i}\cdot\textbf{z}^{text}_{i}/\tau_{2}}}{\sum_{j=1}^{L}\mathrm{e}^{\textbf{z}^{3d}_{i}\cdot\textbf{z}^{text}_{j}/\tau_{2}}}.$ (4)
Here $L$ indicates the number of scenes in a batch, and $\tau_{2}$ is the temperature.
@@ 4.1 Datasets
Our proposed approach is evaluated on two challenging 3D indoor detection datasets: SUN RGB-D, which consists of 5,285 training samples with oriented 3D bounding box labels for 46 object categories. We select the 10 most frequent classes as base classes and the remaining 36 classes as novel classes; ScanNetv2 [3], it has 1,201 training samples with 200 object categories. We use the 10 most frequent classes as base classes and the remaining 50 most frequent classes as novel classes. All configurations above are the same as in [2].
@@ 4.3 Implementation Details
Training Strategy and Hyperparameters. Our training procedure includes three stages. In the first stage, i.e. base detector training, we train the 3D detector for 1000 epochs for a fair comparison with [2]. In the second stage, i.e. novel class discovery, we train for an additional 200 epochs. In the final stage, i.e. feature space alignment, we train another 50 epochs to align the 3D feature space to the vision-language feature space. For the first stage, we configure a batch size of 8 and a learning rate of 0.0004, utilizing 128 queries. Subsequently, during the second and third stages, we specifically adjust the learning rate to 0.0001, batch size to 16, and increase the query size to 196 for enhanced detector training and feature alignment. We set the temperatures $\tau_{1}$ and $\tau_{2}$ to 1.0 and the threshold $\epsilon$ to 0.75. Additionally, the hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are initially set to 0.02 during warm-up and later adjusted to 1, 1, and 0.5, respectively.
Model Selection and Prompt Setting. We modify 3DETR [15] as our 3D detector by removing the classification head and using only the bounding box regression loss. The pretrained Detic [29] is utilized as our 2D detector. For feature alignment, we leverage the pretrained CLIP [20] to encode image and text features. In the class-level and scene-level feature space alignment, we generate text prompts for language feature encoding. Specifically, for the class-level prompt, we generate feature embeddings using the template “A photo of [class name].”, where [class name] represents the category name. For the scene-level prompt, we concatenate the class names in a scene into a list and insert the list into the template “A room with [class list].”, where [class list] denotes the position to insert the list. Additionally, we include the non-object description prompt, i.e., “A photo of nothing.” to represent areas without discriminative objects.
Table 1: Comparison results(%) of our INHA and baseline methods on SUN RGB-D. The label “Need image” signifies that image input is required for inference.
@@ 4.4 Main Results
Method toilet bed chair sofa dresser table cabinet bookshelf pillow sink Avg.
L3DET [31] 56.34 36.15 16.12 23.02 8.13 23.12 14.73 17.27 23.44 27.94 24.63
OV-3DET [13] 57.29 42.26 27.06 31.5 8.21 14.17 2.98 5.56 23 31.6 24.36
CoDA [2] 68.09 44.04 28.72 44.57 3.41 20.23 5.32 0.03 27.95 45.26 28.76